a machine is producing parts that on average a 0.5 probability of being defective. if a random sample of 5 parts is selected and each part is tested, find the fallowing probabilities
a) getting exactly 1 defective part?
b)getting no defective part?
c) getting at most 1 defective part?
d) getting atleast 1 defective part e) getting 5 defective parts?

Answer by stanbon(75887)

a machine is producing parts that on average a 0.5 probability of being defective. if a random sample of 5 parts is selected and each part is tested, find the fallowing probabilities
---
Probability Problems with n = 5 and p = 1/2:
-----------------------
a) getting exactly 1 defective part?
P(x = 1) = 5C1*(1/2)(1/2)4 = 5/32
--------------------------------------------
b)getting no defective part?
P(x = 0) = 5C0(1/2)^0(1/2)^5 = (1/2)^5 = 1/32
--------------------------------------------
c) getting at most 1 defective part?
P(x = 0 or 1) = 1/32 + 5/32 = 6/32
---------------------------------------------
d) getting atleast 1 defective part e) getting 5 defective parts?
P(1<= x <=5) = 1 - P(x = 0) = 1-(1/32) = 31/32
================================================
Cheers,
Stan H.
============